# Continual Domain Randomization

### v1

@@ II Related Work
Continual learning (CL) algorithms enable knowledge accumulation from a sequence of tasks without catastrophic forgetting [5]. CL has traditionally targeted supervised learning scenarios using static datasets or reinforcement learning problems with discrete actions [7, 8]. More recently, CL has been applied to more challenging scenarios such as robot learning, both in simulation [20, 21] as well as in the real world [22, 23]. Continual reinforcement learning (CRL) in continuous action and observation spaces [24] has also emerged as a popular means of training robots to learn multiple tasks. CRL scenarios can be of different types [24]: (i) the state transition and action space for different tasks remain the same, but the reward function changes (the agent needs to perform fundamentally different actions such as pushing, reaching, etc., for each task) [25]; (ii) the state transition function changes, but the reward function and action space remain the same. Here, the environment dynamics change, but the ultimate goal of the agent remains the same [26]. The sim2real scenario belongs to the latter. Progressive networks [27] have also been used for continual sim2real transfer [28]. For instance, as shown by Rusu et al., [28], an initial neural network can be trained in simulation to solve a task, and a new network, with lateral connections to the previous network for feature reusing, is instantiated and trained on the real robot. Another approach used for sim2real transfer is Policy distillation [29]. For instance, Traoré et al., [30] trained a separate policy and generated a dataset of observation-action pairs for each task in simulation. The datasets are then used in a supervised learning fashion to train a single model that can solve the tasks in the real environment. Unlike other domain randomization approaches, in CDR no fixed set of randomization parameters during training is assumed – the model can train on a subset of randomization parameters at a time and adapt to new randomization parameters later. This is achieved by combining sequential randomization with continual learning, which retains the effects of previous randomizations and avoids overfitting to the last set of parameters. Unlike other continual sim2real approaches, CDR uses a single neural network trained across all tasks. Additionally, the online CDR variant avoids the linear increase in memory for keeping a separate dataset or network parameters per task. Finally, CDR can easily be combined with methods like automated [14] or active [15] DR to find suitable ranges for the subset of randomization parameters within a single training run in the sequence.
@@ III Methodology
Figure 1: Overview of our proposed CDR approach. CDR-$\lambda$ (blue arrow) uses a set of network snapshots and Fisher matrices (one for each past task) for continual learning, while CDR-$\mathcal{O}\lambda$ (green arrow) uses only a single parameter snapshot and Fisher matrix. Other baselines are shown with gray arrows.
@@ III-B Continual Domain Randomization
$\mathcal{L}(\theta)=\mathcal{L}_{C}(\theta)+\lambda_{B}||\theta-\theta^{*}_{B}||^{2}_{F_{B}}+\lambda_{A}||\theta-\theta^{*}_{A}||^{2}_{F_{A}}$ (1)
@@
$\mathcal{L}_{i}(\theta)=\tilde{\mathcal{L}}_{i}(\theta)+\cfrac{1}{2}||\theta-\theta_{i-1}^{*}||^{2}_{\gamma F^{*}_{i-1}}$ (3)
@@ IV-B Grasper Task
$\mathbf{a}_{t}=[\dot{q}_{1}^{\prime},\dot{q}_{2}^{\prime},\dot{q}_{3}^{\prime},\dot{q}_{4}^{\prime},\dot{q}_{5}^{\prime},\dot{q}_{6}^{\prime},\dot{q}_{7}^{\prime},]\in\mathbb{R}^{7}$ (9)
@@ IV-E Evaluation Procedure and Metrics
Continuity cost ($\mathcal{C}$[40]: A measure of the smoothness of motion. $\mathcal{C}=\mathbb{E}_{t}\left[\left(\cfrac{a_{t+1}-a_{t}}{\Delta^{a}_{\mathrm{max}}}\right)^{2}\right]$ (12) where $a_{t+1}$ and $a_{t}$ are consecutive actions, and $\Delta^{a}_{\mathrm{max}}$ denotes the maximum difference between consecutive actions during the evaluation rollout. $\mathcal{C}$ ranges between 0 (constant actions) to 1 (each consecutive action reaches opposing limits in the action space). Lower values denote a smoother robot motion and are desirable in the real system.
@@ V-A Reacher
The evaluation results in a non-randomized simulation and on the real system for each randomization strategy for the reaching task at the end of training are presented in Tab. III. The Ideal model performs best in simulation. However, it overfits to the simulation, and this leads to poor transfer to the real system, resulting in lower rewards ($r_{\mathrm{ep}}$), high continuity cost ($\mathcal{C}$), and unstable jittering motions around the target which results in high average distance to the target ($d_{\mathrm{tgt}}$). Randomized has lower rewards in simulation but achieves very good sim2real transfer, leading to higher rewards, low continuity cost, and lowest distance to target on the real system. The CDR models also achieve very good sim2real transfer and perform comparably to or better than the fully Randomized baseline. Due to the stronger EWC regularization during training, CDR-$\lambda$ reaches lower rewards and higher distance to target w.r.t. the CDR-$\mathcal{O}\lambda$ strategy, while in has comparable continuity cost to CDR-$\mathcal{O}\lambda$. Overall, CDR-$\mathcal{O}\lambda$ achieves the best rewards on the real system and has distance and continuity cost values similar to the fully Randomized baseline. Finetuning performs comparably to Ideal in simulation but better on the real system. However, due to the lack of any regularization, it finds over-confident policies in simulation, leading to much worse sim2real transfer than the CDR strategies. The qualitative differences in performance between the different strategies are also noticeable in the supplementary video.
@@ V-C Effect of Continual Learning Regularization Strength
In the context of regularization-based CL methods used to implement CDR, the regularization strength $\lambda$ plays a key role. A very low value makes the CDR models behave like Finetuning, while a very high value makes it difficult to adapt to new randomizations after being initially trained in ideal simulation. We perform additional experiments to investigate the effect of the regularization strength during continual learning of the different randomizations by our CDR models. For this study, we employ the same training procedure as in the reaching experiment in Sec. V-A. We use the same hyperparameters (Tab. I) and only change the value of the EWC $\lambda$ parameter, testing a range of different values of $\lambda=x^{y}$ with $x\in\{1,5\}$ and $y\in\{0,\cdots,4\}$. We train models of CDR-$\lambda$ and CDR-$\mathcal{O}\lambda$ in the ideal simulation followed by two different randomization sequences $TLN$ and $NLT$. We repeated them three times with independent seeds. The evaluation results of the models in the ideal simulation and on the real system are shown in Fig. 6.
@@ VI Discussion and Conclusion
Our experimental results show that full randomization reduces the possibility of finding a good policy in simulation compared to a non-randomized simulation, and it also takes longer to converge. Nevertheless, randomization is important for transferring the model to the real system. Sequential randomization starting from a model pre-trained on non-randomized simulation offers a middle ground between no randomization and full randomization. However, the ordering of the randomizations can affect sim2real transfer. Unfortunately, in zero-shot transfer (without testing on the target system), the effect of each randomization parameter cannot be known with certainty. This can be detrimental since common sequential randomization strategies overfit to the latest randomization they were trained on.
@@
A limitation of our current training approach on one randomization at a time is that we factorize the space of randomization parameters. The complex interactions between separate randomizations in the sequence may not be captured, which may lead to sub-optimal solutions. A possible remedy for this is to define the randomization tasks as sets of different and possibly intersecting groups of randomization parameters. Furthermore, this paper does not address how to define the ranges for the randomization parameters. Excessive randomization is possible even when a single parameter is randomized at a time, e.g., 10% noise randomization translates to $\pm$ 200mm uncertainty of the box’s position and makes the grasping task challenging to solve, leading to poor grasp success on the real system. To address this issue, CDR could be combined with automated domain randomization [14] or active domain randomization [15] to find suitable ranges for each randomization parameter.

### v2

@@ I Introduction
{textblock}
13.24(1.7,0.3) © 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ II Related Work
Continual learning (CL) algorithms enable knowledge accumulation from a sequence of tasks without catastrophic forgetting [5]. CL has traditionally targeted supervised learning scenarios using static datasets or reinforcement learning problems with discrete actions [7, 8]. More recently, CL has been applied to more challenging scenarios such as robot learning, both in simulation [20] as well as in the real world [21, 22]. Continual reinforcement learning (CRL) in continuous action and observation spaces [23] has also emerged as a popular means of training robots to learn multiple tasks. CRL scenarios can be of different types [23]: (i) the state transition and action space for different tasks remain the same, but the reward function changes (the agent needs to perform fundamentally different actions such as pushing, reaching, etc., for each task) [24]; (ii) the state transition function changes, but the reward function and action space remain the same. Here, the environment dynamics change, but the ultimate goal of the agent remains the same [25]. The sim2real scenario belongs to the latter. Progressive networks [26] have also been used for continual sim2real transfer [27]. For instance, as shown by Rusu et al., [27], an initial neural network can be trained in simulation to solve a task, and a new network, with lateral connections to the previous network for feature reusing, is instantiated and trained on the real robot. Another approach used for sim2real transfer is Policy distillation [28]. For instance, Traoré et al., [29] trained a separate policy and generated a dataset of observation-action pairs for each task in simulation. The datasets are then used in a supervised learning fashion to train a single model that can solve the tasks in the real environment. Unlike other domain randomization approaches, in CDR no fixed set of randomization parameters during training is assumed – the model can train on a subset of randomization parameters at a time and adapt to new randomization parameters later. This is achieved by combining sequential randomization with continual learning, which retains the effects of previous randomizations and avoids overfitting to the last set of parameters. Unlike other continual sim2real approaches, CDR uses a single neural network trained across all tasks. Additionally, the online CDR variant avoids the linear increase in memory for keeping a separate dataset or network parameters per task. Finally, CDR can easily be combined with methods like automated [14] or active [15] DR to find suitable ranges for the subset of randomization parameters within a single training run in the sequence.
@@ III Methodology
[…]
@@ III-B Continual Domain Randomization
$\mathcal{L}(\theta)=\mathcal{L}_{C}(\theta)+\cfrac{\lambda_{B}}{2}||\theta-\theta^{*}_{B}||^{2}_{F_{B}}+\cfrac{\lambda_{A}}{2}||\theta-\theta^{*}_{A}||^{2}_{F_{A}}$ (1)
@@
$\mathcal{L}_{i}(\theta)=\tilde{\mathcal{L}}_{i}(\theta)+\cfrac{\lambda}{2}||\theta-\theta_{i-1}^{*}||^{2}_{\gamma F^{*}_{i-1}}$ (3)
@@ IV-B Grasper Task
$\mathbf{a}_{t}=[\dot{q}_{1}^{\prime},\dot{q}_{2}^{\prime},\dot{q}_{3}^{\prime},\dot{q}_{4}^{\prime},\dot{q}_{5}^{\prime},\dot{q}_{6}^{\prime},\dot{q}_{7}^{\prime}]\in\mathbb{R}^{7}$ (9)
@@ IV-E Evaluation Procedure and Metrics
Continuity cost ($\mathcal{C}$[39]: A measure of the smoothness of motion. $\mathcal{C}=100\times\mathbb{E}_{t}\left[\left(\cfrac{a_{t+1}-a_{t}}{\Delta^{a}_{\mathrm{max}}}\right)^{2}\right]$ (12) where $a_{t+1}$ and $a_{t}$ are consecutive actions, and $\Delta^{a}_{\mathrm{max}}$ denotes the maximum difference between consecutive actions during the evaluation rollout. $\mathcal{C}$ ranges between 0 (constant actions) and 100 (each consecutive action reaches opposing limits in the action space). Lower values denote a smoother robot motion and are desirable in the real system.
@@ V-A Reacher
The evaluation results in a non-randomized simulation and on the real system for each randomization strategy for the reaching task at the end of training are presented in Tab. III. The Ideal model performs best in simulation. However, it overfits to the simulation, and this leads to poor transfer to the real system, resulting in lower rewards ($r_{\mathrm{ep}}$), high continuity cost ($\mathcal{C}$), and unstable jittering motions around the target which results in high average distance to the target ($d_{\mathrm{tgt}}$). Randomized has lower rewards in simulation but achieves very good sim2real transfer, leading to higher rewards, low continuity cost, and lowest distance to target on the real system. The CDR models also achieve very good sim2real transfer and perform comparably to or better than the fully Randomized baseline. Due to the stronger EWC regularization during training, CDR-$\lambda$ reaches lower rewards and higher distance to target w.r.t. the CDR-$\mathcal{O}\lambda$ strategy, while it has comparable continuity cost to CDR-$\mathcal{O}\lambda$. Overall, CDR-$\mathcal{O}\lambda$ achieves the best rewards on the real system and has distance and continuity cost values similar to the fully Randomized baseline. Finetuning performs comparably to Ideal in simulation but better on the real system. However, due to the lack of any regularization, it finds over-confident policies in simulation, leading to much worse sim2real transfer than the CDR strategies. The qualitative differences in performance between the different strategies are also noticeable in the supplementary video.
@@ V-C Effect of Continual Learning Regularization Strength
In the context of regularization-based CL methods used to implement CDR, the regularization strength $\lambda$ plays a key role. A very low value makes the CDR models behave like Finetuning, while a very high value makes it difficult to adapt to new randomizations after being initially trained in ideal simulation. Taking this into account, in our main experiments we have used a value for $\lambda$ solely based on empirical evaluation in simulation in order to maintain the zero-shot sim2real condition. Here, we perform additional experiments to provide better intuition on a suitable range for $\lambda$.
For this study, we employ the same training procedure as in the reaching experiment in Sec. V-A. We use the same hyperparameters (Tab. I) and only change the value of the EWC $\lambda$ parameter, testing a range of different values of $\lambda=x\times 10^{y}$ with $x\in\{1,5\}$ and $y\in\{0,\cdots,4\}$. We train models of CDR-$\lambda$ and CDR-$\mathcal{O}\lambda$ in the ideal simulation followed by two different randomization sequences $TLN$ and $NLT$. We repeated them three times with independent seeds. The evaluation results of the models in the ideal simulation and on the real system are shown in Fig. 6.
@@ VI Discussion and Conclusion
In general, full randomization of many parameters simultaneously makes the task more difficult [14, 15, 4], and our experiments show that it takes longer for the models to converge under this setting. Sequential randomization starting from a model pre-trained on non-randomized simulation offers a middle ground between no randomization and full randomization. However, the ordering of the randomizations can affect sim2real transfer. Unfortunately, in zero-shot transfer (without testing on the target system), the effect of each randomization parameter cannot be known with certainty. This can be detrimental since common sequential randomization strategies overfit to the latest randomization they were trained on.
@@
A limitation of our current training approach on one randomization at a time is that we factorize the space of randomization parameters. The complex interactions between separate randomizations in the sequence may not be captured, which may lead to sub-optimal solutions. A possible remedy for this is to define the randomization tasks as sets of different and possibly intersecting groups of randomization parameters [41]. Furthermore, this paper does not address how to define the ranges for the randomization parameters. Excessive randomization is possible even when a single parameter is randomized at a time, e.g., 10% noise randomization translates to $\pm$ 200mm uncertainty of the box’s position and makes the grasping task challenging to solve, leading to poor grasp success on the real system. To address this issue, CDR could be combined with automated domain randomization [14] or active domain randomization [15] to find suitable ranges for each randomization parameter.